• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name J. Singh

    Log entry time 21:49:30 on July12,2003

    Entry number 103786

    keyword=Septum Training: 401.4A

    started a septum training today at 1629. previous high was 394A. quenched at 2050 @ 401.4A.

    started ramp with 0.110V. @ 200A increased to 0.115V. After 350A, ramp slowed to well below 1A/min. i couldn't change the ramp voltage b/c the Hall was locked due to laser alignment. really slow ramp didn't seem to help that much. next time i will increase the ramp voltage to a much higher value (~0.125V).


    FIGURE 1